Show, by perturbing the geodesic in the equatorial plane, that the circular orbit of a photon at r= = 3m is unstable.
Answer #1


Schwarzschild Solution: Geodesics ř - Zvi? – re”ő? – re“ sin? 062 +22 )t? = 0 (1) = r = 0 = 2 2 Ö+r0 sin 0 cos 002 0 (2) cosThe geodesics on the equatorial plane are: 1 1 ï vj2 12 – reu + e2v jt? = : 0 (6) 2 2 2 + = 0 r = 0 Ï +vrt -e į² tel; 2 +An observer with 4-velocity UP will find that the energy of a particle with 4-momentum pH is: E = Pu UM An observer at rest aBy combining the 2 integrals of motion and eqn (9) we can eliminate the proper time to derive an equation for a 3D path of thTrajectories of photons Photons, as any zero rest mass particle, move on null geodesics. For photons we cannot use the properRadial motion of photons For radial motion 0 = 0 and we get e” į? - e-V;2= 0 Particle Observer ci from which we obtain dr dtCircular motion of photons For circular orbits we have r=constant and thus from eqn (41) we see that the only possible radiusStability of photon orbits We can rewrite the energy equation (40) as eu 1 12 L2 + (44) r2 D2 where D=L/E and Veff = eu/r V